### v1

@@ Appendix C Empirical results
$\beta_{0}$ 1.108*** 0.060** 0.004** 0.060*** 0.004*** -0.067*** (0.008) (0.009)
(0.021) (0.031) (0.002) (0.009) (0.001) (0.025) (0.055) (0.077)
$\alpha_{11}$ 0.995*** 0.167*** 0.153*** 0.167*** 0.153*** -0.415*** 0.143*** 0.150***
(0.162) (0.010) (0.011) (0.016) (0.020) (0.014) (0.016) (0.014)
$\alpha_{12}$ 0.904*** -0.036*** -0.032*** -0.036 -0.032 0.735*** -0.032 -0.022
(0.164) (0.005) (0.006) (0.027) (0.035) (0.019) (0.023) (0.021)
$\alpha_{13}$ 0.084 -0.015** 0.017*** -0.015 0.017 -0.198*** -0.003 -0.026***
(0.158) (0.007) (0.005) (0.010) (0.036) (0.006) (0.008) (0.006)
$\alpha_{14}$0.251 -0.004*** -0.004 -0.004 -0.004 -0.160*** -0.002 -0.019***
(0.160) (0.001) (0.003) (0.005) (0.012) (0.007) (0.005) (0.004)
$\alpha_{21}$ 0.001 0.092*** 0.088*** 0.092*** 0.088*** -0.006*** 0.037*** 0.037***
(0.007) (0.015) (0.016) (0.004) (0.029) (0.001) (0.003) (0.003)
$\alpha_{22}$ 0.001 0.031*** 0.019*** 0.031*** 0.019*** -0.037*** 0.013*** 0.014***
(0.008) (0.012) (0.003) (0.002) (0.003) (0.003) (0.002) (0.002)
$\alpha_{23}$ 0.000 0.004 0.035 0.004*** 0.035*** 0.014*** 0.007*** -0.003***
(0.007) (0.021) (0.022) (0.001) (0.004) (0.001) (0.001) (0.001)
$\alpha_{24}$ 0.005 0.003 0.006 0.003*** 0.006*** 0.003*** 0.002*** 0.002***
(0.008) (0.004) (0.008) (0.000) (0.001) (0.001) (0.000) (0.000)
$\alpha_{31}$ 0.018 0.087 0.068 0.087 0.068 -0.027 0.041 0.041
(1.320) (1.065) (1.012) (0.127) (0.126) (0.863) (0.079) (0.166)
$\alpha_{32}$ 0.020 0.029 0.025 0.029 0.025 -0.081 0.013 0.006
(0.685) (2.392) (2.003) (0.268) (0.212) (0.062) (0.091) (0.099)
$\alpha_{33}$ 0.003 0.028 0.005 0.028 0.005 0.019 0.010 0.009
(0.331) (5.691) (4.001) (0.316) (0.305) (0.891) (1.063) (0.085)
$\alpha_{34}$ 0.005 -0.022 -0.020 -0.022 -0.020 0.019 -0.011 -0.013
(0.369) (0.355) (0.093) (0.287) (0.365) (1.605) (0.062) (0.076)
@@
$\mathscr{T}_{n}$ 0.865 0.397 0.473 0.981 0.486 0.983 0.829 0.600
$\mathscr{T}_{n}^{*}$ 0.873 0.393 0.453 0.953 0.467 0.927 0.833 0.580
$H_{0}:\mu=0$ 86.301 90.233
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@
$n$ 7355
The result in the first column does not consider any spatial weight matrix.
Figure 1: Comparison of parametric method and semiparametric method in Section 6.1, capital.
Figure 2: Comparison of parametric method and semiparametric method in Section 6.1, labor.
Figure 3: Comparison of parametric method and semiparametric method in Section 6.1, CRS.
Table 8: Data descriptions in Section 6.2.
Variables Descriptions Mean Min Mean SD
$y$ log(MEDV) 3.035 1.609 3.912 0.408
$x_{1}$ log(LSTAT) 2.371 0.548 3.637 0.601
$x_{2}$ log(RAD) 1.868 0 3.178 0.875
$p_{1}$ log(CRIM) -0.780 -5.064 4.488 2.162
$p_{2}$ log(RM) 1.832 1.270 2.172 0.112
$p_{3}$ log(TAX) 5.931 5.231 6.567 0.396
$z$ log(DIS) 1.188 0.122 2.495 0.540
Table 9: Estimation and testing results in Section 6.2 using polynomial series with $h=2$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$\beta_{0}$ 4.081*** 3.987*** 4.081*** 3.987*** -0.003 0.132*** 0.022*
(0.267) (0.272) (0.147) (0.156) (0.006) (0.018) (0.012)
$\beta_{1}$0.432*** -0.432*** -0.432*** -0.432*** -0.143* -0.053*** -0.162
(0.025) (0.024) (0.027) (0.025) (0.075) (0.004) (0.175)
$\beta_{2}$ 0.166*** 0.154*** 0.166*** 0.154*** 0.086* 0.025*** 0.044**
(0.024) (0.023) (0.018) (0.017) (0.049) (0.001) (0.020)
$\alpha_{11}$0.044*** -0.041*** -0.044*** -0.041* -0.056* -0.033** -0.017***
(0.006) (0.006) (0.006) (0.024) (0.030) (0.001) (0.004)
$\alpha_{12}$ 0.028*** 0.025*** 0.028*** 0.025*** 0.043* 0.022*** 0.014***
(0.004) (0.004) (0.005) (0.005) (0.022) (0.001) (0.003)
$\alpha_{21}$ 0.042** 0.046** 0.042** 0.046*** 0.115** -0.001 -0.068***
(0.019) (0.018) (0.019) (0.006) (0.051) (0.002) (0.011)
$\alpha_{22}$0.012 -0.015 -0.012 -0.015 0.136*** 0.008*** 0.035***
(0.012) (0.011) (0.009) (0.022) (0.049) (0.001) (0.004)
$\alpha_{31}$$0**$027*** -$0**$027*** $0*$005*** -0.019***
(0.007) (0.007) (0.006) (0.002) (0.030) (0.002) (0.003)
$\alpha_{32}$ 0.076* 0.070* 0.076*** $0**$084*** 0.082*** 0.101***
(0.040) (0.040) (0.003) (0.004) (0.015) (0.021) (0.011)
@@
$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 183.122 91.459 123.439
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
$H_{0}:\mu=0$ 32.336 35.141
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.013
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 10: Estimation and testing results in Section 6.2 using trigonometric series with $h=2$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
(0.049) (0.044) (0.028) (0.023) (0.069) (0.027) (0.057)
$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
(0.149) (0.040) (0.083)
$\beta_{0}$ 3.527*** 3.547*** 3.527*** 3.547*** 0.001 0.567*** 0.675***
(0.236) (0.229) (0.175) (0.139) (0.009) (0.091) (0.028)
$\beta_{1}$0.410*** -0.415*** -0.410*** -0.415*** -0.235*** -0.396*** -0.456***
(0.024) (0.023) (0.023) (0.024) (0.064) (0.021) (0.032)
$\beta_{2}$ 0.156*** 0.146*** 0.156*** 0.146*** 0.170*** 0.154*** 0.170***
(0.024) (0.023) (0.017) (0.017) (0.031) (0.018) (0.020)
$\alpha_{11}$0.056* -0.035* -0.056*** -0.035*** -0.024* -0.006** -0.007*
(0.040) (0.021) (0.003) (0.003) (0.012) (0.003) (0.004)
$\alpha_{12}$0.023*** -0.022*** -0.023*** -0.022*** -0.163*** -$0**$024***
(0.005) (0.005) (0.006) (0.005) (0.027) (0.005) (0.005)
$\alpha_{21}$ 0.054*** 0.049*** 0.054*** 0.049*** 0.241** 0.037*** 0.048***
(0.010) (0.010) (0.008) (0.007) (0.097) (0.009) (0.011)
$\alpha_{22}$0.016 -0.007 -0.016*** -0.007*** 0.212*** 0.106** 0.132**
(0.018) (0.018) (0.003) (0.002) (0.026) (0.052) (0.065)
$\alpha_{31}$0.027*** -0.023*** -0.027*** -0.023*** -0.017 -0.025*** -0.030***
(0.004) (0.004) (0.002) (0.002) (0.036) (0.002) (0.002)
$\alpha_{32}$ 0.010* 0.009* 0.010*** 0.009** 0.622*** 0.053*** 0.065***
(0.006) (0.005) (0.004) (0.004) (0.015) (0.004) (0.005)
$H_{0}:\alpha=0$ 38.897 34.423 229.115 325.327 266.301 175.332 144.982
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.020 0.007 0.007 0.020 0.007 0.013 0.007
$H_{0}:\mu=0$ 28.883 21.326
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 11: Estimation and testing results in Section 6.2 using polynomial series with $h=4$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$\beta_{0}$ 3.285*** 2.925*** 3.285 2.925*** -$0**$092* -0.440**
(0.464) (0.444) (7.140) (0.347) (0.003) (0.047) (0.211)
$\beta_{1}$0.399*** -0.386*** -0.399*** -0.386*** -0.516*** 0.021* -0.105***
(0.026) (0.025) (0.039) (0.025) (0.092) (0.012) (0.013)
$\beta_{2}$ 0.158*** 0.137*** 0.158*** 0.137*** 0.313*** 0.019* 0.052***
(0.025) (0.025) (0.023) (0.018) (0.057) (0.011) (0.013)
$\alpha_{11}$ 0.041 0.031 0.041 0.031 0.048* 0.034** 0.006**
(0.053) (0.050) (0.067) (0.039) (0.027) (0.015) (0.003)
$\alpha_{12}$0.224* -0.185* -0.224* -0.185** -0.120*** -0.072** -0.012*
(0.118) (0.112) (0.135) (0.086) (0.024) (0.034) (0.006)
$\alpha_{13}$ 0.228*** 0.193** 0.228*** 0.193*** -0.055* 0.050*** 0.080*
(0.087) (0.083) (0.087) (0.062) (0.028) (0.016) (0.046)
$\alpha_{14}$0.064*** -0.055*** -0.064*** -0.055*** 0.058*** -0.011** -0.018*
(0.021) (0.020) (0.018) (0.015) (0.018) (0.005) (0.010)
$\alpha_{21}$0.706*** -$0**$706*** -$0**$073 0.019 0.003**
(0.240) (0.226) (0.226) (0.114) (0.073) (0.011) (0.001)
$\alpha_{22}$ 1.574*** 1.596*** 1.574*** 1.596*** -0.014 -0.026* -0.006**
(0.484) (0.453) (0.460) (0.229) (0.045) (0.015) (0.002)
$\alpha_{23}$1.040*** -1.078*** -1.040*** -1.078*** -0.020 0.056*** 0.030***
(0.312) (0.291) (0.295) (0.147) (0.016) (0.014) (0.009)
$\alpha_{24}$ 0.217*** 0.230*** 0.217*** 0.230*** -0.057* $0**$045**
(0.065) (0.060) (0.060) (0.030) (0.030) (0.003) (0.018)
$\alpha_{31}$ 0.262*** 0.264*** 0.262*** 0.264*** 0.068*** -$0**$017***
(0.084) (0.079) (0.080) (0.041) (0.015) (0.003) (0.006)
$\alpha_{32}$0.572*** -0.577*** -0.572*** -0.577*** 0.032 0.019*** 0.035**
(0.166) (0.155) (0.164) (0.079) (0.027) (0.007) (0.014)
$\alpha_{33}$ 0.371*** 0.381*** 0.371*** 0.381*** -0.026 -0.089* -0.021***
(0.106) (0.099) (0.105) (0.050) (0.035) (0.046) (0.004)
$\alpha_{34}$0.076*** -$0**$076*** -0.080*** 0.029* 0.011* 0.038***
(0.022) (0.020) (0.021) (0.010) (0.015) (0.006) (0.007)
@@
$H_{0}:\alpha=0$ 33.876 20.486 392.020 928.892 601.967 394.776 352.386
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.013 0.013 0.007 0.020 0.013 0.013
$H_{0}:\mu=0$ 56.124 66.981
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 12: Estimation and testing results in Section 6.2 using trigonometric series with $h=4$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$\beta_{0}$ 3.207*** 2.762*** 3.207*** 2.762*** -0.005 0.001** 0.001***
@@
$\beta_{1}$0.392*** -0.376*** -0.392*** -0.376*** -0.296*** -0.257** -0.200***
(0.026) (0.024) (0.031) (0.024) (0.061) (0.111) (0.025)
$\beta_{2}$ 0.146*** 0.124*** 0.146*** 0.124*** -0.046 0.103 0.069***
(0.025) (0.024) (0.019) (0.016) (0.036) (0.299) (0.010)
$\alpha_{11}$0.089*** -0.082*** -0.089*** -0.082*** 0.234*** -0.074** -0.060***
(0.027) (0.025) (0.019) (0.021) (0.078) (0.032) (0.017)
$\alpha_{12}$ 0.201*** 0.175*** 0.201*** 0.175*** 0.355*** 0.162** 0.068*
(0.068) (0.064) (0.046) (0.045) (0.093) (0.071) (0.038)
$\alpha_{13}$$0**$084*** -$0**$084*** -0.315*** -0.075** -0.020
(0.031) (0.029) (0.022) (0.019) (0.095) (0.033) (0.017)
$\alpha_{14}$$0**$081*** -$0**$081*** 0.236*** -0.073*** -0.054***
(0.026) (0.025) (0.019) (0.021) (0.066) (0.022) (0.017)
$\alpha_{21}$ 0.086*** 0.104*** 0.086*** 0.104*** -0.283*** 0.064*** 0.107***
(0.031) (0.029) (0.021) (0.016) (0.085) (0.019) (0.011)
$\alpha_{22}$0.611*** -0.688*** -0.611*** -0.688*** -0.005 -0.394 -0.506***
(0.179) (0.165) (0.170) (0.146) (0.062) (1.827) (0.138)
$\alpha_{23}$ 0.354*** 0.396*** 0.354*** 0.396*** 0.305*** 0.229*** 0.278***
(0.103) (0.095) (0.100) (0.086) (0.108) (0.071) (0.080)
$\alpha_{24}$ 0.025 0.050 0.025 $0**$042 0.036 0.091***
(0.036) (0.034) (0.019) (0.014) (0.033) (0.076) (0.011)
$\alpha_{31}$0.027** -0.027** -0.027*** -0.027*** 0.128*** -0.025*** -0.035***
(0.013) (0.012) (0.009) (0.007) (0.032) (0.006) (0.005)
$\alpha_{32}$ 0.216*** 0.238*** 0.216*** 0.238*** -0.138 0.150* 0.217***
(0.061) (0.056) (0.057) (0.048) (0.086) (0.088) (0.043)
$\alpha_{33}$0.123*** -0.136*** -0.123*** -0.136*** -0.047 -0.086** -0.121***
(0.035) (0.033) (0.033) (0.029) (0.046) (0.038) (0.025)
$\alpha_{34}$0.004 -0.012 -0.004 -0.012** -0.014 -0.009 -0.030***
(0.013) (0.012) (0.008) (0.005) (0.026) (0.018) (0.005)
@@
$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 258.704 91.459 123.439
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
$H_{0}:\mu=0$ 43.155 46.397
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@ Appendix D Simulation results
Table 14: Empirical sizes for $\mathbb{W}_{1}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
@@
7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.072 0.102 0.038 0.072 0.094 0.048 0.076 0.092
$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.088 0.029 0.069 0.118 0.019 0.078 0.127
$h=4$ $\mathscr{T}_{n}$ 0.044 0.090 0.118 0.032 0.080 0.110 0.036 0.082 0.110
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.088 0.078 0.127 0.137 0.019 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.058 0.132 0.184 0.032 0.106 0.146 0.038 0.064 0.094
$\mathscr{T}_{n}^{*}$ 0.049 0.088 0.137 0.028 0.127 0.137 0.019 0.098 0.108
7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.100 0.132 0.026 0.076 0.110 0.048 0.082 0.110
$\mathscr{T}_{n}^{*}$ 0.029 0.137 0.167 0.020 0.088 0.098 0.020 0.059 0.118
$h=4$ $\mathscr{T}_{n}$ 0.048 0.100 0.142 0.024 0.066 0.100 0.026 0.066 0.100
$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.186 0.019 0.078 0.118 0.019 0.078 0.137
$h=8$ $\mathscr{T}_{n}$ 0.060 0.120 0.178 0.042 0.074 0.120 0.018 0.054 0.104
$\mathscr{T}_{n}^{*}$ 0.038 0.176 0.196 0.039 0.078 0.147 0.020 0.078 0.137
7.1 $h=2$ $\mathscr{T}_{n}$ 0.062 0.108 0.140 0.028 0.058 0.092 0.026 0.056 0.098
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.008 0.020 0.039 0.019 0.088 0.118
$h=4$ $\mathscr{T}_{n}$ 0.072 0.116 0.170 0.020 0.066 0.116 0.036 0.078 0.108
$\mathscr{T}_{n}^{*}$ 0.049 0.078 0.127 0.029 0.069 0.088 0.015 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.094 0.168 0.216 0.026 0.062 0.112 0.034 0.078 0.136
$\mathscr{T}_{n}^{*}$ 0.058 0.127 0.186 0.029 0.108 0.167 0.014 0.088 0.137
trigonometric series $n=100$ $n=300$ $n=500$
@@
7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.064 0.098 0.044 0.070 0.104 0.046 0.070 0.092
$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.078 0.029 0.039 0.088 0.019 0.048 0.118
$h=4$ $\mathscr{T}_{n}$ 0.052 0.092 0.122 0.030 0.076 0.140 0.032 0.074 0.116
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.069 0.039 0.108 0.118 0.019 0.058 0.137
$h=8$ $\mathscr{T}_{n}$ 0.068 0.138 0.180 0.038 0.110 0.134 0.034 0.066 0.094
$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.137 0.029 0.118 0.157 0.019 0.038 0.108
7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.084 0.112 0.038 0.084 0.118 0.048 0.084 0.114
$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.176 0.029 0.059 0.108 0.010 0.039 0.088
$h=4$ $\mathscr{T}_{n}$ 0.050 0.108 0.152 0.022 0.066 0.104 0.032 0.070 0.114
$\mathscr{T}_{n}^{*}$ 0.069 0.118 0.186 0.049 0.098 0.127 0.019 0.058 0.127
$h=8$ $\mathscr{T}_{n}$ 0.064 0.108 0.162 0.032 0.086 0.124 0.022 0.052 0.080
$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.020 0.108 0.157 0.019 0.058 0.157
7.1 $h=2$ $\mathscr{T}_{n}$ 0.060 0.090 0.120 0.034 0.070 0.112 0.022 0.062 0.098
$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.098 0.020 0.039 0.078 0.018 0.068 0.137
$h=4$ $\mathscr{T}_{n}$ 0.064 0.094 0.164 0.022 0.076 0.114 0.032 0.070 0.106
$\mathscr{T}_{n}^{*}$ 0.078 0.098 0.147 0.039 0.088 0.098 0.017 0.068 0.118
$h=8$ $\mathscr{T}_{n}$ 0.098 0.158 0.206 0.024 0.060 0.102 0.032 0.078 0.120
$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.039 0.127 0.167 0.016 0.068 0.127
Table 15: Empirical sizes for $\mathbb{W}_{2}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$

### v2

@@ Appendix C Empirical Results
$H_{0}:\mu=0$ 86.301*** 90.233***
Standard errors in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@
(a) Capital
(b) Labor
(c) CRS
Figure 1: Cobb-Douglas versus semiparametric production.
Table 3: Estimation and testing results in Section 6.2
Panel A: polynomial series $h=2$
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$H_{0}:\alpha=0$ 40.041*** 33.913*** 290.165*** 33.913*** 183.122*** 91.459*** 123.439***
$H_{0}:\mu=0$ 32.336*** 35.141***
Panel B: polynomial series $h=4$
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$H_{0}:\alpha=0$ 33.876*** 20.486*** 392.020*** 928.892*** 601.967*** 394.776*** 352.386***
$H_{0}:\mu=0$ 56.124*** 66.981***
Panel C: trigonometric series $h=2$
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$H_{0}:\alpha=0$ 40.041*** 33.913*** 290.165*** 33.913*** 258.704*** 91.459*** 123.439***
$H_{0}:\mu=0$ 43.155*** 46.397***
Panel D: trigonometric series $h=4$
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
@@
$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
(0.149) (0.040) (0.083)
@@
$H_{0}:\alpha=0$ 38.897*** 34.423*** 229.115*** 325.327*** 266.301*** 175.332*** 144.982***
$H_{0}:\mu=0$ 28.883*** 21.326***
Panel E: Comparison of parametric method and semiparametric method in Section 6.2
@@ Appendix D Simulation Results
Table 4: Empirical sizes for $\mathbb{W}_{1}$, rejection probabilities at 1%, 5% and 10% levels.
$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
@@
7.1 $h=2$ asy-p 0.035 0.075 0.115 0.033 0.063 0.098 0.042 0.074 0.107
chi-p 0.013 0.051 0.112 0.006 0.049 0.096 0.010 0.056 0.106
$h=4$ asy-p 0.033 0.074 0.112 0.031 0.074 0.107 0.038 0.077 0.114
chi-p 0.013 0.056 0.106 0.017 0.050 0.105 0.010 0.055 0.106
$h=8$ asy-p 0.033 0.073 0.117 0.025 0.060 0.091 0.011 0.060 0.096
chi-p 0.014 0.057 0.108 0.011 0.046 0.088 0.008 0.040 0.091
7.1 $h=2$ asy-p 0.036 0.077 0.111 0.049 0.086 0.109 0.034 0.071 0.089
chi-p 0.015 0.048 0.109 0.014 0.062 0.107 0.007 0.044 0.087
$h=4$ asy-p 0.038 0.090 0.126 0.052 0.092 0.125 0.027 0.070 0.094
chi-p 0.015 0.059 0.118 0.015 0.070 0.117 0.005 0.045 0.094
$h=8$ asy-p 0.029 0.069 0.101 0.023 0.072 0.105 0.030 0.075 0.106
chi-p 0.012 0.055 0.097 0.009 0.053 0.097 0.013 0.054 0.102
7.1 $h=2$ asy-p 0.042 0.072 0.111 0.037 0.085 0.113 0.025 0.059 0.091
chi-p 0.012 0.057 0.106 0.011 0.059 0.112 0.004 0.039 0.088
$h=4$ asy-p 0.038 0.082 0.125 0.031 0.082 0.116 0.025 0.069 0.102
chi-p 0.013 0.056 0.114 0.012 0.058 0.111 0.007 0.046 0.096
$h=8$ asy-p 0.023 0.066 0.108 0.031 0.067 0.098 0.027 0.069 0.104
chi-p 0.011 0.043 0.100 0.012 0.048 0.095 0.011 0.052 0.103
$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
@@
7.1 $h=2$ asy-p 0.037 0.084 0.123 0.037 0.071 0.101 0.033 0.081 0.111
chi-p 0.012 0.055 0.120 0.011 0.049 0.098 0.008 0.060 0.106
$h=4$ asy-p 0.040 0.086 0.141 0.034 0.072 0.101 0.030 0.073 0.112
chi-p 0.011 0.058 0.131 0.011 0.052 0.099 0.014 0.046 0.104
$h=8$ asy-p 0.020 0.053 0.086 0.021 0.056 0.089 0.024 0.045 0.090
chi-p 0.009 0.042 0.083 0.008 0.038 0.081 0.008 0.035 0.088
7.1 $h=2$ asy-p 0.048 0.083 0.110 0.045 0.083 0.115 0.036 0.069 0.100
chi-p 0.016 0.061 0.108 0.015 0.062 0.113 0.014 0.055 0.099
$h=4$ asy-p 0.049 0.093 0.129 0.046 0.095 0.125 0.035 0.084 0.117
chi-p 0.015 0.070 0.121 0.017 0.070 0.119 0.012 0.051 0.115
$h=8$ asy-p 0.031 0.066 0.106 0.020 0.058 0.086 0.016 0.041 0.086
chi-p 0.013 0.049 0.103 0.005 0.042 0.085 0.003 0.033 0.086
7.1 $h=2$ asy-p 0.037 0.074 0.107 0.048 0.078 0.111 0.038 0.080 0.103
chi-p 0.016 0.049 0.106 0.014 0.060 0.107 0.010 0.055 0.101
$h=4$ asy-p 0.039 0.089 0.113 0.035 0.071 0.107 0.032 0.071 0.110
chi-p 0.014 0.069 0.110 0.015 0.057 0.099 0.011 0.047 0.103
$h=8$ asy-p 0.033 0.063 0.096 0.019 0.037 0.069 0.016 0.052 0.088
chi-p 0.015 0.050 0.090 0.007 0.032 0.062 0.006 0.032 0.088
Table 5: Empirical sizes for $\mathbb{W}_{2}$, rejection probabilities at 1%, 5% and 10% levels.
$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$